### v2

@@ 2.2 Shape generalization of the proposed method
Figure 5: Analaysis of shape generalization of the proposed method. a. The proposed method reconstructs the object field where the diffraction intensities are synthesized by incorporating distinct amplitude and phase pairs. b. The proposed method trained on diffraction intensity does not meet expectations to reconstruct the object fields. c. Estimated likelihood steers sampling trajectory toward conditioned distribution $p({\bm{x}}|{\bm{y}})$.
We further conducted an in-depth exploration of the proposed method through simulation studies. Specifically, we synthesized diffraction intensities using the angular spectrum method, with amplitude and phase maps intentionally decorrelated in morphology. In this process, we used the amplitude image of polystyrene beads in conjunction with two distinct phase counterparts: one derived from a tissue section and the other from a natural image. Remarkably, the proposed method demonstrates its exceptional shape generalization capability (by restoring complex amplitude accurately), not requiring the inherent similarity between amplitude and phase components (Figure 5a). To discern whether the generalization performance predominantly stems from the conditioning methods, we conducted an ablation experiment in which the diffusion prior was omitted from the sampling procedure or replaced by inaccurate alternatives. The results reveal a deviation of the generated object field from the learned data manifold, leading to a notable reconstruction failure. Additionally, to establish a boundary for shape generalization, we trained the diffusion model using defocused amplitudes of polystyrene beads (i.e., out-of-focus) and subsequently applied it to reconstruct object fields from synthetic measurements. In Figure 5b, it is evident that the diffusion prior, trained on diffracted patterns, does not achieve the same level of generalization as observed in Figure 5a. This indicates that the shape generalization boundary encompasses a range of focused object shapes. These findings highlight the pivotal role of the diffusion prior in solving ill-posed inverse problems by enabling superior shape generalization through test-time optimization within the context of posterior sampling.

### v3

@@ 2.2 Shape generalization of the proposed method
Figure 5: Analaysis of shape generalization of the proposed method. a. The proposed method reconstructs the object field where the diffraction intensities are synthesized by incorporating distinct amplitude and phase pairs. b. Ablation Study: Comparison of the reconstructed complex amplitude corresponding to the synthesized measurement. The reconstruction is shown for: (Top) A randomly initialized diffusion model (i.e., without a diffusion prior), and (Bottom) Applying the Tweedie’s formula during the reconstruction procedure to estimate the clean phase image $\hat{x}_{0}^{p}$. c. Visualization of the sampling trajectory for the proposed method and the two ablation cases detailed in b (w/o diffusion prior and w/ tweedie for phase).
We further conducted an in-depth exploration of the proposed method through simulation studies. Specifically, we synthesized diffraction intensities using the angular spectrum method, with amplitude and phase maps intentionally decorrelated in morphology. In this process, we used the amplitude image of polystyrene beads in conjunction with two distinct phase counterparts: one derived from a tissue section and the other from a natural image. Remarkably, the proposed method demonstrates its exceptional shape generalization capability (by restoring complex amplitude accurately), not requiring the inherent similarity between amplitude and phase components (Figure 5a). To discern whether the generalization performance predominantly stems from the conditioning methods, we conducted an ablation experiment in which the diffusion prior was omitted from the sampling procedure. The results reveal a deviation of the generated object field from the learned data manifold, leading to a notable reconstruction failure as shown in Figure 5b.
Additionally, when an inaccurate phase estimation is used by incorporating a diffusion model trained solely on amplitude data, similar reconstruction failures occur (Figure 5b). Interestingly, in both cases, we observed that the reconstructions correspond to realizations of ill-posed solutions. In other words, the reconstructed field reproduces the measurement when the propagation distance matches that of the observation (Figure 5c). However, when the propagation distance is altered, the reconstructed complex field fails to yield a physically meaningful diffraction pattern. This behavior provides strong evidence of the problem’s ill-posed nature and highlights the pivotal role of the diffusion prior. These findings demonstrate the proposed method’s effectiveness by enabling superior shape generalization through test-time optimization within the context of posterior sampling.